Montserrat Volcano Observatory, Montserrat, West Indies

Scientific Report 58
22 March 1996

This report covers the 7-day period to 24:00 on the 22nd March.


Seismicity has remained at a generally low level. Low intensity swarms of volcano-tectonic and hybrid events have been recorded. These are generally longer in duration than in previous weeks. Dome growth has continued in the summit area of the January 20 dome and much of this material is now spreading out over the upper surface of the October dome. Pyroclastic flows have again occurred over the Galway's Wall during earthquake swarms.

Visual Observations

Visibility during the early part of the week was poor. However toward the latter part of the week the volcano was almost completely clear allowing excellent chances for observations of the changes to the dome complex. Growth has continued in the uppermost areas of the January 20 dome which is now advancing over the October dome. Incandescence has been observed from new spines and from material exposed during rockfalls at night.

The eastern and south-eastern faces of the dome continue to be active, producing rockfalls and small to moderate pyroclastic flows. The largest of these flows originated from the upper south-eastern area of the January 20 dome and ran along the southern crater wall into the upper part of the Tar River Valley. Rapid changes of the summit area of the January 20 dome occurred between the 19th and 21st when spines collapsed to be replaced by new spines. The overall appearance of the southern part of the January 20 dome is now more blocky. A period of ash venting was observed from the lower part of the eastern face of the dome on March 21st, lasting for approximately 30 minutes.

Pyroclastic flows have continued to occur over the Galway's Wall. These have frequently been related to large hybrid events which have occurred during earthquake swarms. Small landslides have also occurred from areas to the east and west of the low point on the crater rim over which the flows travel. An erosional chute has been formed in the pyroclastic flow deposit leading out from the crater wall. Fresh deposits seen on March 18th which had formed overnight had well developed levee structures. The run-out distance of these flows is gradually increasing and they have now reached over 1 km from the crater wall.

New fractures were observed on March 20th running through the southern buttress of Galway's Wall, in the area adjacent to Perches Mountain. These fractures are between 100-150 metres in length and trend 1700.


Table 1: Earthquake types
These earthquake counts are of events that triggered the short-period seismic network event recording system between 0000 and 2400 each day.

Date		VT	LP	Hybrid	Dome RF
16 MAR 97	1	0	6	6
17 MAR 97	1	2	159	10
18 MAR 97	4	0	52	2
19 MAR 97	29	0	88	24
20 MAR 97	0	1	12	15
21 MAR 97	8	3	24	19
22 MAR 97	37	4	50	22

Table 2: VT / hybrid earthquake swarms.

Date		Start time	Durn (hrs)	VTs	Hybrid
17 MAR 97	06:44		37.85		5	211
19 MAR 97	08:29		25.28		29	97
21 MAR 97	06:37		5.27		7	12
22 MAR 97	12:09		15.35		36	51

The level of seismic activity during the week has been generally low, and about the same level as previous weeks. Low-intensity swarms of volcano-tectonic and hybrid earthquakes continued, with the swarms lasting for longer than in previous weeks. The character of the swarms has also changed- previously the swarms started with impulsive (volcano-tectonic) earthquakes of various magnitudes, which graded into small-magnitude, emergent "hybrid" earthquakes. The swarms during this period had both types of events during the entire swarm. In reality the VT and hybrid classifications are end members of a continuous suite of earthquake types.

Ground Deformation

EDM measurements were made on the northern triangle (Windy Hill-Farrells-St. George's Hill) and on the lines from Lower Amersham-Upper Amersham and Lower Amersham-Chance's Steps. The slant distance of the lines to the Farrells target site are erratic but an overall stable trend continues. The Lower-Upper Amersham slant distance is stable but the Lower Amersham-Chance's Step appears to be showing a very slight shortening trend. There is also a suggestion of slow subsidence of the Upper Amersham and Chance's Steps target sites.

A 20-hour GPS occupation of the Harris and Hermitage stations was made on the 17th-18th March. The Hermitage station has not shown any more movement since March 11th.

Dome Volume Measurements

A survey was completed on 19 March using a combination of the fixed location photographic method and the GPS/range-finding binocular method. A GPS survey of the talus at the base of the dome was undertaken on 18 March and combined with the data from the survey of 19 March. A volume of 0.8 x 106m3 has been added to the dome since 14 March, giving an average extrusion rate of 1.67m3/s during this 5 day period.

Growth of the dome continues to be concentrated in the summit region and from the photographic profiles it is apparent that the region of active growth is moving westward towards the back of the dome and is now over spilling the trace of the January scar. This region of the dome is only partially visible from the photographic location points so this latest extrusion rate is considered to be a minimum value. From the profiles the summit region has grown by 15m in the period from 14 - 19 March. Growth maps from the surveys also indicated a significant amount of talus material accumulating in the southern moat between the dome and eastern portion of the Galway's wall. This accumulation is due to material being shed from the top of the dome on the southern side down a well developed chute. A second chute has developed on the NE from material being shed from the northern portion of the summit this is causing slow but continuous erosion of the October material.

An important development that has become apparent during this last survey is that the Pre-September scar material which surrounds the dome on the north-western, western and south-western sides is being pushed outwards by the growing dome. The amount of movement is difficult to constrain without further GPS surveys but small amounts of movement can now be traced as far back as November. Displacements listed below are outward movements of the top of the Pre-September scar as measured from photographs from Whites.

Date				Displacement	Rate 
23 November - 8 January		3.9m		0.08m/day
8 January - 19 March		9.1m		0.13m/day

As the rate of displacement appears to be increasing and as the dome within the Pre-September scar continues to grow it would not be unexpected for rockfall activity on the outside, western edge of the scar to be reinitiated in the near future.

A GPS - bathymetry survey of the water depth around the pyroclastic fan at the foot of the Tar River Valley was made on 21st March. Over 500 data points were collected around the fan to depths of 63m, allowing better definition of the submarine topography. When combined with the survey of the fan surface on 12th February, a total fan volume of 15.5 x 106 m3 was derived.

Environmental Monitoring

Measurements of sulphur dioxide flux were made using the MiniCOSPEC on 17 March (317 t/d). This is about the same level as in the previous week.

Rain water samples were collected at 4 locations around the volcano on 16 March, and results are shown in Table 3. One sample was collected from the overflow of Trials reservoir in Fairfield.

These results show that the rainwater directly west of the volcano continues to be highly acidic and has high concentrations of certain anions. The sample from Trials is within World Health Organisation levels for all analytes.

Table 3: Rain and surface water geochemistry, 16 Mar 97
Units: conductivity (mS/cm), total dissolved solids (g/l), sulphate (mg/l), chloride (mg/l), fluoride (mg/l), nd (not detected)

Location		pH 	Cond.	TDS	Slphts	Chlrds	Flrds
Upper Amersham		2.39	1.883	0.942	34	232	1.35
Lower Amersham		2.70	0.731	0.366	8	100	>1.5
Police HQ, Plymouth	2.81	0.571	0.285	3	68	1.4
Weekes			6.11	0.070	0.035	nd	14.4	0.1
Trials reservoir	7.55	0.659	0.330	40	83	0.55

Ash was collected at several sites to the west of the volcano. No further results on ash analyses were received during this period. Dust sampling was re-started at 5 sites in and around Plymouth on 17th March. Samples will be sent back to the UK for analysis of particle size and mineral content.


MVO offers it thanks to Inspector Morson and his crew for the use of the police launch on March 21st for the bathymetric survey and their help throughout the day.

Staff Changes
Desmond Seupersad (SRU)

Richard Luckett (BGS)

Montserrat Volcano Observatory